Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

light bulb fails

Expert replies
by sanju09 » Mon Mar 12, 2012 2:28 am
A string of 10 light bulbs is wired in such a way that if any individual light bulb fails, the entire string fails. If for each individual light bulb the probability of failing during time period T id 0.06, what is the probability that the string of light bulbs will fail during time period T?
A.0.06
B.(0.06)^10
C.1-(0.06)^10
D.(0.94)^10
E.1-(0.94)^10
The mind is everything. What you think you become. –Lord Buddha



Sanjeev K Saxena
Quantitative Instructor
The Princeton Review - Manya Abroad
Lucknow-226001

www.manyagroup.com
Join the discussion
Source: — Problem Solving |

by killer1387 » Mon Mar 12, 2012 2:38 am
required probability= 1- probability that all 10 bulbs dont fail
= 1- (1-0.06)^10
= 1- 0.94^10

IMO E
Join the discussion

by Anurag@Gurome » Mon Mar 12, 2012 5:12 am
sanju09 wrote:A string of 10 light bulbs is wired in such a way that if any individual light bulb fails, the entire string fails. If for each individual light bulb the probability of failing during time period T id 0.06, what is the probability that the string of light bulbs will fail during time period T?
A.0.06
B.(0.06)^10
C.1-(0.06)^10
D.(0.94)^10
E.1-(0.94)^10
Probability of 1 light bulb failing during time period T = 0.06
So, probability of 1 light bulb NOT failing during time period T = 1 - 0.06 = 0.94

For the string to be successful, all the light bulbs should pass because if any individual light bulb fails, the entire string fails.
Probability of 10 light bulbs NOT failing = (0.94)^10
Probability of 10 light bulbs failing = 1 - Probability of 10 light bulbs NOT failing = 1 - (0.94)^10

The correct answer is E.
Anurag Mairal, Ph.D., MBA
GMAT Expert, Admissions and Career Guidance
Gurome, Inc.
1-800-566-4043 (USA)

Join Our Facebook Groups
GMAT with Gurome
https://www.facebook.com/groups/272466352793633/
Admissions with Gurome
https://www.facebook.com/groups/461459690536574/
Career Advising with Gurome
https://www.facebook.com/groups/360435787349781/
Join the discussion

by kul512 » Mon Mar 12, 2012 1:34 pm
Just adding some information-

One thing which must be taken care is that failing of one bulb is not going to affect the failing of other bulb. means these events are mutually exclusive events and that's why we need to use
"1- none fails" formula rather than "1st fails OR 2nd fails OR 3rd fails....".

Means we cants simply add all the probability.
Sometimes there is very fine line between right and wrong: perspective.
Join the discussion